# Stabilizing MoE Reinforcement Learning by Aligning Training and Inference Routers

### v1

@@ Abstract
Yudong Wang†‡ Zhifang Sui†⋄ Fuli Luo
@@ 1 Introduction
A critical challenge in LLM-based RL is balancing efficiency and stability, with the latter being essential for reliable performance. Modern RL frameworks typically employ distinct engines for inference and training phases (e.g., SGLang (Zheng et al., 2024) for rollout and Megatron (Shoeybi et al., 2019) for training). This architectural separation can lead to divergent token probabilities, potentially causing catastrophic RL collapse (He and Lab, 2025). To mitigate this discrepancy, Yao et al. (2025) incorporate importance sampling mechanism in policy updating, while He and Lab (2025) introduce specialized compute kernels to reduce nondeterminism during LLM Inference. However, in practice, existing approaches do not fully resolve the intensified off-policy issue that arises during RL training on Mixture-of-Experts (MoE) models.
@@
On realistic RLVR tasks with MoE models, R3 demonstrates significant superiority in training stability, and performance. Our method consistently shows marked improvements in both efficiency and overall performance when compared against existing approaches designed to stabilize RL training. Furthermore, its applicability to both on-policy and mini-batch style off-policy RL scenarios underscores the robustness of our approach.
@@
    We proposed Rollout Routing Replay, which reuses inference-time routing distributions inside the training engine to align routing behavior between training and inference.
@@ 2 Preliminaries
A critical inconsistency arises from the common practice of using separate engines for rollout and training, where data is sampled via an inference policy ($\pi_{\text{infer}}$) but the loss is computed using a training policy ($\pi_{\text{train}}$), as shown in Equation 1. This policy mismatch drives training instability in reinforcement learning, an issue that, within MoE models, we identify as stemming primarily from router inconsistency. The proposed solution alleviates this fundamental issue, making it broadly applicable, orthogonal to, and compatible with recent policy optimization frameworks like GRPO (Shao et al., 2024), GSPO (Zheng et al., 2025) and DAPO (Yu et al., 2025).
@@ 3 Training-Inference Discrepancies
Training-inference discrepancies in RL frameworks frequently lead to unstable training and model collapse. In this section, we demonstrate that this policy mismatch is significantly amplified in MoE models, primarily stemming from inconsistent routing distributions Furthermore, we observe that even multiple runs of the same training framework can produce divergent token probabilities, further contributing to RL training instability.
@@ 4.3 Empirical Analysis of R3 on Training-Inference Discrepancies
To evaluate the effectiveness of R3 in reducing training-inference discrepancies, we repeat the procedure described in Section 3.1 with Qwen3-30B-A3B model. In this process, we cache the routing distributions obtained during inference on SGLang and replay them within the Megatron framework. After obtaining the inference engine probability and training engine probability, using the equation 2, we estimate the KL divergence between training and inference. Results shows that after applying R3, the KL divergence between training and inference decrease from $1.5\times 10^{-3}$ to $7.5\times 10^{-4}$, which is near the $6.4\times 10^{-4}$ observed for the dense model. This intuitively indicates a reduction in the training-inference discrepancy. We also drew the cumulative distribution plot of the ratio of training-inference discrepancies in Fig. 2(d) with R3. The plot indicates that, for the MoE model, applying R3 reduces by an order of magnitude the frequency of tokens with large training-inference discrepancies.
@@ 5.1 Setting
Task and Dataset We choose mathematical reasoning as the target task for training. For the training dataset, we collect and filter problems from many open-source datasets, including BigMath (Albalak et al., 2025), ORZ (Hu et al., 2025), and others, yielding approximately 100,000 verifiable math problems. For the evaluation dataset, we adopt AIME24, AIME25, AMC23, and MATH500 (level 5) as our benchmark datasets. We report Avg@32 for AIME24 and AIME25, Avg@16 for AMC23, and Avg@4 for MATH500 (level 5). During training, some models may experience performance degradation or even collapse in later stages. To ensure a fair evaluation, we measure model performance every 5 global steps throughout a single training run. We then report the highest observed performance along with the corresponding training step at which it occurs.
@@
Baseline Methods We consider the following baseline optimization methods for comparison: (a) GRPO (Shao et al., 2024), additionally applies the Clip Higher technique from DAPO (Yu et al., 2025), with parameters $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.27$; (b) TIS (Yao et al., 2025), uses an upper clipping threshold $C=2$; (c) GSPO (Zheng et al., 2025), employs sequence-level importance sampling, with parameters $\epsilon_{\text{low}}=3\times 10^{-4}$ and $\epsilon_{\text{high}}=4\times 10^{-4}$. Since our R3 method is orthogonal to optimizers such as GSPO or TIS, we also evaluate various combinations of these techniques.
Multiple Mini Steps vs. Single Mini Step In PPO-like algorithms, a batch of samples collected in one global step is typically divided into multiple mini steps so that the policy can be updated several times. However, previous research has shown that applying a strictly on-policy strategy with only a single mini step can yield better performance. We investigate both settings. For multiple mini steps, we set the mini step count to 8, meaning each PPO mini step trains $2048/8=256$ samples with 8 optimizer updates and a learning rate of $1\times 10^{-6}$. In the multiple mini-step scenario, for R3 method we replayed the routing both when recomputing the old policy and updating policy. For the single mini step scenario, we set the mini step count to 1, meaning all 2048 samples are updated once. Due to fewer updates, we increase the learning rate to $3\times 10^{-6}$. We don’t recompute old policy in this scenario.
@@
Other Setting We implement R3 using the VeRL (Sheng et al., 2024) framework, use Megatron (Shoeybi et al., 2019) for training, and SGLang (Zheng et al., 2024) for inference. We set the batch size to 256 and $n=8$, totaling 2048 samples per round. The maximum prompt length is 2048, and the maximum generation length is 30720. We adopt the Dynamic Sampling strategy from Yu et al. (2025), retaining only partially correct samples during generation until a sufficient batch size is accumulated for training. No auxiliary loss for expert balancing is introduced.
@@ 5.2 Experimental Results and Analysis
The main evaluation results are shown in Table 1. More detailed evaluation results and training logs, including per-evaluation scores as well as training metrics such as entropy, average reward, gradient norm, and training-inference KL divergence, are provided in the appendix A and B
@@ 6.2 Instability in Reinforcement Learning for LLMs
Reinforcement Learning with verifiable rewards (RLVR) has become a standard method for refining the complex reasoning (Xie et al., 2025), mathematical (Shao et al., 2024; Hu et al., 2025), and code abilities (Luo et al., 2025b) of large language models. Algorithms such as GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025) are now widely adopted. However, this training paradigm also faces challenges with training instability, which can hinder convergence and risk model collapse (Yang et al., 2025; Xia et al., 2025; Chen et al., 2025). Recent research has identified several root causes and proposed corresponding solutions. For instance, Zhao et al. (2025b) aim to stabilize training by optimizing the geometric mean of token-level rewards. Separately, Zheng et al. (2025) posit that GRPO’s instability stems from a misapplication of importance sampling weights and introduce a new sequence-level importance ratio to correct it.
Yao et al. (2025) propose Truncated Importance Sampling (TIS) to mitigate the inconsistency stemming from discrepancies between training frameworks (e.g., FSDP (Zhao et al., 2023) and inference engines (e.g., vLLM (Kwon et al., 2023)). Alternatively, He and Lab (2025) attribute instability to non-determinism in compute kernels, proposing batch-invariant operations, though this incurs significant performance overhead and has not been explored for RL on MoE models. These stability challenges are critically exacerbated in Mixture-of-Experts (MoE) architectures. We find this heightened instability is primarily caused by routing inconsistencies: the router’s sensitivity, combined with the framework gap, causes expert assignments for the same response to differ between the rollout and training phases. To directly address this core issue, our proposed method, Rollout Routing Replay, explicitly records and replays the inference-time routing decisions during the training pass. This simple strategy enforces consistency, enabling MoE models to achieve training stability comparable to their dense counterparts without incurring computational overhead.
@@ 7 Conclusion
Appendix A Detail Evaluation Results
Appendix B Detail Training Metrics

### v2

@@ Abstract
Yudong Wang†‡ Zhifang Sui†⋄ Fuli Luo§⋄
@@
00footnotetext: Work done during internship at Xiaomi Corporation.00footnotetext: Co-corresponding authors.
@@ 1 Introduction
A critical challenge in LLM-based RL is balancing efficiency and stability, with the latter being essential for reliable performance. Modern RL frameworks typically employ distinct engines for inference and training phases (e.g., SGLang (Zheng et al., 2024) for rollout and Megatron (Shoeybi et al., 2019) for training). This architectural separation can lead to divergent token probabilities, potentially causing catastrophic RL collapse (T. M. Lab, 2025). To mitigate this discrepancy, Yao et al. (2025) incorporate an importance-sampling mechanism in policy updating, while T. M. Lab (2025) introduce specialized compute kernels to reduce nondeterminism during LLM inference. However, in practice, existing approaches do not fully resolve the intensified off-policy issue that arises during RL training on Mixture-of-Experts (MoE) models.
@@
On realistic RLVR tasks with MoE models, R3 demonstrates significant improvements in training stability and performance. Our method consistently shows marked improvements in both efficiency and overall performance when compared against existing approaches designed to stabilize RL training. Furthermore, its applicability to both on-policy and mini-batch style off-policy RL scenarios underscores the robustness of our approach.
@@
    We propose Rollout Routing Replay, which reuses inference-time routing distributions inside the training engine to align routing behavior between training and inference.
@@ 2 Preliminaries
A critical inconsistency arises from the common practice of using separate engines for rollout and training, where data is sampled via an inference policy ($\pi_{\text{infer}}$) but the loss is computed using a training policy ($\pi_{\text{train}}$), as shown in Equation 1. This policy mismatch causes training instability in reinforcement learning. In MoE models, we find that it mainly stems from router inconsistency. Our proposed solution effectively mitigates this issue, making it broadly applicable, orthogonal to, and compatible with recent policy optimization frameworks such as GRPO (Shao et al., 2024), GSPO (Zheng et al., 2025), and DAPO (Yu et al., 2025).
@@ 3 Training-Inference Discrepancies
Training-inference discrepancies in RL frameworks frequently lead to unstable training and model collapse. In this section, we demonstrate that this policy mismatch is significantly amplified in MoE models, primarily stemming from inconsistent routing distributions. Furthermore, we observe that even multiple runs of the same training framework can produce divergent token probabilities, further contributing to RL training instability.
@@ 4.1 Implementation
Through meticulous system design and implementation, we enable the storage and retrieval of MoE router mask data from the inference engine, with an overall latency overhead below 3% during the rollout phase.
@@ 4.3 Empirical Analysis of R3 on Training-Inference Discrepancies
To evaluate the effectiveness of R3 in reducing training-inference discrepancies, we repeat the procedure described in Section 3.1 with the Qwen3-30B-A3B model. In this process, we cache the routing distributions obtained during inference on SGLang and replay them within the Megatron framework. After obtaining the inference engine probability and training engine probability, using the equation 2, we estimate the KL divergence between training and inference. The result shows that after applying R3, the KL divergence between training and inference decrease from $1.5\times 10^{-3}$ to $7.5\times 10^{-4}$, which is near the $6.4\times 10^{-4}$ observed for the dense model. This intuitively indicates a reduction in the training-inference discrepancy. We also draw the cumulative distribution plot of the ratio of training-inference discrepancies in Fig. 2(d) with R3. The plot indicates that, for the MoE model, applying R3 reduces by an order of magnitude the frequency of tokens with large training-inference discrepancies.
@@ 5.1 Setting
Task and Dataset We choose mathematical reasoning as the target task for training. For the training dataset, we collect and filter problems from many open-source datasets, including Big-Math-RL-Verified [^1] https://huggingface.co/datasets/SynthLabsAI/Big-Math-RL-Verified (Albalak et al., 2025), ORZ-Math [^2] https://huggingface.co/Open-Reasoner-Zero/datasets (Hu et al., 2025), and others, yielding approximately 100,000 verifiable math problems. For the evaluation dataset, we adopt AIME24, AIME25, AMC23, and MATH500 (level 5)(Hendrycks et al., 2021) as our benchmark datasets. We report Avg@32 for AIME24 and AIME25, Avg@16 for AMC23, and Avg@4 for MATH500 (level 5). During training, some models may experience performance degradation or even collapse in later stages. To ensure a fair evaluation, we evaluate the model performance every 5 global steps and report the highest observed performance along with the corresponding training step at which it occurs.
@@
Methods For our R3 method, we cache the routing during rollout and replay it both when recomputing the old policy and updating policy. We consider the following baseline optimization methods for comparison: (a) GRPO (Shao et al., 2024), additionally applies the Clip Higher technique from DAPO (Yu et al., 2025), with parameters $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.27$; (b) TIS (Yao et al., 2025), uses an upper clipping threshold $C=2$; (c) GSPO (Zheng et al., 2025), employs sequence-level importance sampling, with parameters $\epsilon_{\text{low}}=3\times 10^{-4}$ and $\epsilon_{\text{high}}=4\times 10^{-4}$. Since our R3 method is orthogonal to optimization techniques such as GSPO or TIS, we also evaluate various combinations of these techniques.
Multiple Mini Steps vs. Single Mini Step In PPO-like algorithms, a batch of samples collected in one global step is typically divided into multiple mini steps so that the policy can be updated several times. However, recent research (Hao et al., 2025; Yue et al., 2025; Ji et al., 2025) has shown that applying only one single mini step can yield better performance. We investigate both settings. For multiple mini steps, we set the mini step to 8 and the learning rate to $1\times 10^{-6}$. For the single mini step scenario, we set the mini step to 1. Due to fewer updates, we increase the learning rate to $3\times 10^{-6}$.
@@
Other Setting We implement R3 using the VeRL (Sheng et al., 2024) framework, use Megatron (Shoeybi et al., 2019) for training, and SGLang (Zheng et al., 2024) for inference. We set the batch size to 256 and $n=8$, totaling 2048 samples per global step. The maximum prompt length is 2048, and the maximum generation length is 30720. We adopt the Dynamic Sampling strategy from Yu et al. (2025), retaining only partially correct samples during generation until a sufficient batch size is accumulated for training. No auxiliary loss for expert balancing is introduced.
@@ 5.2 Experimental Results and Analysis
The main evaluation results are shown in Table 1. More detailed evaluation results and training logs, including per-evaluation scores as well as training metrics such as entropy, average reward, gradient norm, and training-inference KL divergence, are provided in the appendix A
@@
5.3 Multi-Turn Task Reinforcement Learning
Qwen3-30B-A3B, mini_step=1, max_global_step=180
Method SWE-bench Verified()($\uparrow$) Best Global Step Crash Step
GRPO 31.80 70 90
GRPO+R3 38.60 160 -
Table 2: Evaluation results of RL training on SWE Task with Qwen3-30B-A3B
In Section 4.2, we mentioned that the R3 method can be adapted to Multi-Turn tasks. To verify this capability, we conduct reinforcement learning experiments on the software engineering (Jimenez et al., 2024) task using the Qwen3-30B-A3B (Yang et al., 2025) model. We use R2E-Gym-Lite [^3] https://huggingface.co/datasets/R2E-Gym/R2E-Gym-Lite (Jain et al., 2025) as the training dataset and SWE-bench Verified [^4] https://huggingface.co/datasets/princeton-nlp/SWE-bench_Verified (Jimenez et al., 2024) as the validation dataset. For training, We set the maximum sequence length to $65536$ tokens (including the prompt, model response, and environment observation), the batch size to $64$, the maximum number of interaction steps to $50$, the learning rate to $2\times 10^{-6}$, and the mini_step to $1$. For evaluation, We report Pass@1 on SWE-bench Verified. We evaluate the model performance every 10 global steps. We compare the performance of GRPO and GRPO+R3 under these settings. All other hyperparameters follow the configuration described in Section 5.1.
The results are presented in Table 2, and the detailed training metrics are provided in Appendix C. The training process of GRPO without R3 collapses after around 90 training steps, while GRPO+R3 maintains stable. The final performance of GRPO+R3 reaches a Pass@1 of 38.6, which is 6.8 points higher than that of GRPO. Since we use the Router Mask Caching technique described in Section 4.2, we do not need to re-prefill the prefix to obtain the routing mask, thus the rollout speed is not slowed down. These results confirm that R3 generalizes well to multi-turn reinforcement learning settings.
@@ 6.2 Instability in Reinforcement Learning for LLMs
Reinforcement Learning with verifiable rewards (RLVR) has become a standard method for refining the complex reasoning (Xie et al., 2025), mathematical (Shao et al., 2024; Hu et al., 2025), and code abilities (Luo et al., 2025b) of large language models. Algorithms such as GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025) are now widely adopted. However, this training paradigm also faces challenges with training instability, which can hinder convergence and risk model collapse (Yang et al., 2025; Xia et al., 2025; Chen et al., 2025). Recent research has identified several root causes and proposed corresponding solutions. For instance, Zhao et al. (2025b) aim to stabilize training by optimizing the geometric mean of token-level rewards. Separately, GSPO (Zheng et al., 2025) posits that GRPO’s instability stems from a misapplication of importance sampling weights and introduce a new sequence-level importance ratio to correct it. Yao et al. (2025) propose Truncated Importance Sampling (TIS) to mitigate the inconsistency stemming from discrepancies between training frameworks (e.g., FSDP (Zhao et al., 2023) and inference engines (e.g., vLLM (Kwon et al., 2023)). Alternatively, (T. M. Lab, 2025) attributes instability to non-determinism in compute kernels, proposing batch-invariant operations, though this incurs significant performance overhead and has not been explored for RL on MoE models. These stability challenges are critically exacerbated in Mixture-of-Experts (MoE) architectures. We find this heightened instability is primarily caused by routing inconsistencies: the router’s sensitivity, combined with the framework gap, causes expert assignments for the same response to differ between the rollout and training phases. To directly address this core issue, our proposed method, Rollout Routing Replay, explicitly records and replays the inference-time routing decisions during the training pass. This simple strategy enforces consistency, enabling MoE models to achieve training stability comparable to their dense counterparts without incurring computational overhead.
6.3 Routing Replay
Routing Replay was first introduced in Zheng et al. (2025) (hereafter referred to as Recompute Routing Replay). We clarify how our Rollout Routing Replay differs from it. As shown in Figure 1, there are three forward-pass stages in a typical reinforcement learning framework: Rollout, Recompute and Update. Recompute Routing Replay caches the routing from the Recompute stage and replays it in the Update stage, addressing the issue of routing discrepancies caused by model updates. However, this approach does not take into account that inconsistencies between training and inference frameworks can also lead to routing discrepancies. In contrast, Our Rollout Routing Replay caches the routing from the Rollout stage and replays it in both the Recompute and Update stages, thereby resolving routing discrepancies arising from framework difference. Notably, since our method uses the same routing for both the Recompute and Update stages, it also addresses routing discrepancies caused by model updates. Furthermore, when the mini step is set to 1, the current policy and the old policy become identical, which makes the Recompute Routing Replay method ineffective. Since discrepancies between training and inference frameworks continue to exist, the Rollout Routing Replay method remains effective under this setting.
@@ 7 Conclusion
Appendix A Detailed Evaluation Results and Training Metrics
Figure 7: The detailed evaluation results of the experiment in Section 5.
Figure 8: The detailed training metrics of the experiment in Section 5.
Appendix B Additional Experiments of Reasoning SFT Model
To evaluate the performance of R3 on reasoning SFT models, we fine-tune Qwen3-30B-A3B-Base (Yang et al., 2025) on the open-source long reasoning dataset Mixture-of-Thoughts  [^5] https://huggingface.co/datasets/open-r1/Mixture-of-Thoughts(Face, 2025; Penedo et al., 2025; Lozhkov et al., 2025; Bercovich et al., 2025), producing the Qwen3-30B-A3B-ReasoningSFT model. We conduct reinforcement learning experiments on this model. We compare GSPO and GRPO+R3 with a learning rate of $2\times 10^{-6}$ when $\text{mini\_step}=4$, and compare GRPO and GRPO+R3 with a learning rate of $3\times 10^{-6}$ when $\text{mini\_step}=1$; all other hyperparameters follow the configuration described in Section 5.1.
The results are shown in Table 3, and training and validation metrics are illustrated in Fig 9. The experiments show that training processes without R3 collapse (for $\text{mini\_step}=4$, GSPO collapses at training step $90$, and for $\text{mini\_step}=1$, GRPO collapses at training step $40$), while training processes with R3 remain robust. In addition, training with R3 exhibits a more stable pattern of sequence-length growth as well as more stable entropy and gradient norms throughout the training process.
Best Metric(Best Global Step)
Qwen3-30B-A3B-ReasoningSFT, mini_step=4, max_global_step=180
GSPO 67.50(35) 48.33(20) 94.21(60) 91.79(35) 74.62(55) 90
GRPO+R3 69.06(165) 51.66(155) 95.15(160) 93.28(145) 77.00(165) -
Qwen3-30B-A3B-ReasoningSFT, mini_step=1, max_global_step=180
GRPO 68.54(25) 48.95(25) 93.90(30) 91.23(20) 75.30(25) 40
GRPO+R3 69.47(85) 53.12(180) 95.46(170) 93.28(180) 76.79(180) -
Table 3: Evaluation results of the RL training process of Qwen3-30B-A3B-ReasoningSFT
Figure 9: The detailed training metrics and evaluation results of the RL training process of Qwen3-30B-A3B-ReasoningSFT
Appendix C Detailed Training Metrics of Multi-Turn Reinforcement Learning
Figure 10: The detailed training metrics and evaluation results of RL training process on SWE Task with Qwen3-30B-A3B